A sound wave from a siren has an intensity of 106.1 W/m2 at a certain point, and a second sound wave from a nearby ambulance has an intensity level 11 dB greater than the siren's sound wave at the same point. What is the intensity level of the sound wave due to the ambulance?

Answer in dB
Answer #1

Intensity level due to first ambulance, beta1 = 10*log(I/Io)

= 10*log(106.1/10^-12)

= 140.26 dB


intensity level deu to second ambulance, beta2 = 140.26 + 11

= 151.26 dB

let I2 is the intensity of sound of second ambulance.

beta2 = 10*log(I2/Io)

151.26 = 10*log(I2/10^-12)

15.126 = log(I2/10^-12)

==> I2 = 10^(15.126-12)

= 1336.6 W/m^2 <<<<<<<<<<<------------Answer